OpenStudy (anonymous):

Simplify (w8 )-6 ___ (w2) Please Helpp

OpenStudy (mrcoolguy):

w8-6

OpenStudy (calculusxy):

\[\large \frac{ w^8 - 6 }{ w^2 } \] \[\large w^8 \div w^2 = w^{8 -2} = w^6\] \[\large w^6 - 6\]

OpenStudy (anonymous):

(w^8/w^2)^-6

OpenStudy (anonymous):

Answers are

OpenStudy (calculusxy):

is the question \[(w^8 \div w^2)^{-6}\]

OpenStudy (anonymous):

a)-w^36 b)-w^24 c)1/w^36 d)1/w^24

OpenStudy (anonymous):

w^8 is on top of w^2 and they are both inside one parentheses and to the power of -6

OpenStudy (anonymous):

plleaase help

OpenStudy (calculusxy):

\((w^6)^{-6}\) we have the power rule \((x^a)^b\) = \(x^{a \times b}\) \(w^{6 \times -6} = w^{-36}\) another rule comes up \(\large x^{-a} = \frac{1}{x^a}\) \(\huge w^{-36} = \frac{1}{w^{36}}\)

OpenStudy (anonymous):

Thank you so muchh god bless you

OpenStudy (calculusxy):

you're welcome :)